OpenStudy (anonymous):

Simplify the expression. 57 – 12 ÷ 4 • 3 Ive got 48

OpenStudy (anonymous):

12 / 4 = 3 3 x 3 = 9 57 - 9 =48

OpenStudy (anonymous):

@jdoe0001 @iambatman @Godlovesme @campbell_st @silverr Am I correct?

OpenStudy (anonymous):

@Loser66 @Compassionate

OpenStudy (anonymous):

@sammixboo @pooja195 @nincompoop

OpenStudy (anonymous):

@dtan5457 @robtobey @lerodge857 @Preetha @poopsiedoodle

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@jdoe0001

OpenStudy (texaschic101):

57 - 12/4 * 3 = 57 - 3 * 3 = 57 - 9 = 48 I agree with you...I got 48

OpenStudy (anonymous):

@texaschic101 Thank you so much! Ill tell you what I get.

OpenStudy (anonymous):

56

OpenStudy (anonymous):

@lerodge857 How'd you do it?

OpenStudy (anonymous):

PEMDAS

OpenStudy (anonymous):

I got it mastered!

OpenStudy (texaschic101):

you can do multiplication/division first...starting from left to right...then do add/subtract, starting from left to right....ending up with 48

OpenStudy (anonymous):

first multiply then divide then subtract!!

OpenStudy (anonymous):

It was 48 It showed me the correct answer and it was 48 for some reason it wasnt PEMDAS It was algebra practice or something like that.

OpenStudy (anonymous):

hmmm I would have done PEMDAS

OpenStudy (anonymous):

Me too But its always incorrect for some reasons!?!

OpenStudy (texaschic101):

pemdas can get confusing....because multiply/divide can be done from left to right in that order...then add/subtract from left to right

OpenStudy (anonymous):

Oh yea thats why!

OpenStudy (anonymous):

well beats me!!!

OpenStudy (texaschic101):

just remember this for your future problems.. P = parenthesis E = exponent M/D -- multiplication/division from left to right -- in that order A/S -- add/subtract from left to right in that order

OpenStudy (anonymous):

so it either or?

OpenStudy (anonymous):

Yea is it!

OpenStudy (texaschic101):

yes...its either or example : 2 + 4/2 * 6 = 2 + 2 * 6 2 + 12 14 division came first, then multiplication, then addition

OpenStudy (texaschic101):

if it would have been : 2 * 4/2 + 6 2 * 2 + 6 4 + 6 10

OpenStudy (texaschic101):

all I am saying is that multiplication/division can be done at the same time if you go from left to right

OpenStudy (texaschic101):

just do all multiplication/division before adding/subtracting

OpenStudy (anonymous):

Thanks :)

OpenStudy (texaschic101):

sure thing :)